OpenStudy (anonymous):

Given: 3x - 2 ≤ 2x + 1. Choose the solution set. {x | x E R, x ≤ -3} {x | x E R, x ≤ -1} {x | x E R, x ≤ 1} {x | x E R, x ≤ 3}

OpenStudy (anonymous):

i added two to all numbers is tht correct so far and i got 3x≤4x+3

OpenStudy (ivettef365):

ok you need to combine like terms so switch equation from one side to another, when you switch remember to change sign as follows: 3x -2x <= 1 + 2 do you see what I did ?

OpenStudy (anonymous):

you moved like terms to one side?

OpenStudy (ivettef365):

right I moved the positive 2x to the left side so I changed to negative 2x and I moved the negative 2 to the other side so I changed to positive 2

OpenStudy (anonymous):

so now do i add a certain number or?

OpenStudy (ivettef365):

now just add you have 3x - 2x = x and you have 1 + 2 = 3 so you have x <= 3

OpenStudy (anonymous):

one sec brb gunna put this on paper...Xp

OpenStudy (anonymous):

so @ivettef365 d

OpenStudy (ivettef365):

correct

OpenStudy (anonymous):

thanks for help

OpenStudy (ivettef365):

yw :)